OpenStudy (lgbasallote):

a car is traveling at 50.0 mi/h on a horizontal highway. if the coefficient of static friction between road and tires on a rainy day is 0.100, what is the minimum dstance in which the car will stop? how do i approach this problem?

OpenStudy (ujjwal):

Retarding force, F=\(\mu\)N or, ma = \(\mu\) mg or, a=\(\mu\)g=0.100 \(\times\) 9.8=0.98 mi/h^2 So, the retarding acceleration is 0.98 mi/h^2 Now, \(v^2=u^2-2as\) .....(-ve sign to indicate that the acceleration is retarding) Final velocity v is zero, since the car finally comes to rest. so, 0=15^2 -2*0.98* s s=110.25 miles.. Hence, the car will stop in 110.25 miles

OpenStudy (lgbasallote):

can you explain what you did @ujjwal ? i mean step-by step...it's pretty hard when you cram everything

OpenStudy (anonymous):

first of all, friction force is being applied on the tires because of which the car is decelerating. then as explained by ujjwal, calculate the deceleration of the car. now substitute that value in v^2 = u^2 + 2as to find out s, which is what u need.

OpenStudy (lgbasallote):

wait...we do FBD first right?

OpenStudy (anonymous):

we don't need an FBD in everything...this sum was quite straightforward...and u don't need to resolve any forces. so u don't need to draw FBD.

OpenStudy (lgbasallote):

F=μN what does that mean?

OpenStudy (anonymous):

that's the formula for frictional force. u = coefficient of friction and N = normal force.

OpenStudy (anonymous):

remember one thing. we use FBD to figure out the net force acting on a body. its a vector diagram.

OpenStudy (lgbasallote):

oh...is this the same as \[\large f_k = \mu_k F_N?\]

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

so this is kinetic friction?

OpenStudy (lgbasallote):

oh wait it's static...

OpenStudy (anonymous):

that's right.

OpenStudy (anonymous):

no...its kinetic.

OpenStudy (lgbasallote):

but the formula for static friction is \[\large f_s \le \mu_s F_N\] how does he know it's eqaul and not less? and the problem says static friction...

OpenStudy (anonymous):

@ujjwal shouldn't it be kinetic? since the tires are in motion?

OpenStudy (anonymous):

@ujjwal got it. "An example of static friction is the force that prevents a car wheel from slipping as it rolls on the ground. Even though the wheel is in motion, the patch of the tire in contact with the ground is stationary relative to the ground, so it is static rather than kinetic friction." got it from WIKI.

OpenStudy (ujjwal):

Lost my internet for a while.. Static friction will work here since the wheels are rolling. They are not sliding!

OpenStudy (anonymous):

what about @Igbasallote's question?

OpenStudy (lgbasallote):

yeah...

OpenStudy (anonymous):

i guess this answers it: The maximum possible friction force between two surfaces before sliding begins is the product of the coefficient of static friction and the normal force: .

OpenStudy (lgbasallote):

uhmm layman's terms?

OpenStudy (anonymous):

and in this case, the tires are not sliding. hence static friction.

OpenStudy (anonymous):

u know in about pure rolling?

OpenStudy (ujjwal):

The frictional force will be less than \(\mu_s\)N as long as the body is not in motion.

OpenStudy (lgbasallote):

well...it is said in the problem that it is static hehe so how do you know if you use \[f_s = \mu_s F_N\] or \[f_s < \mu_s F_N\]

OpenStudy (ujjwal):

Once the body is in motion, F=\(\mu_s\)N

OpenStudy (anonymous):

yea....that explains it.

OpenStudy (lgbasallote):

that applies always when in motion?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

so what's the next step? it's determined that \[f_s = \mu_s F_N\] now what?

OpenStudy (anonymous):

but from Newton's 2nd law, F = ma.....sub this in place of frictional force. then proceed how ujjwal showed b4.

OpenStudy (lgbasallote):

\[f_s = F = ma???\]

OpenStudy (anonymous):

yes

OpenStudy (ujjwal):

@lgbasallote Answer to your above question: No. That applies in case of motion where the body is rolling and you can use \(\mu_s\)N for frictional force. Else you have to use coefficient of kinetic friction.

OpenStudy (lgbasallote):

but mass and acceleration are not given...how will that help?

OpenStudy (ujjwal):

you are supposed to find acceleration and mass cancels out as i have done it.

OpenStudy (anonymous):

@Igbasallote: whenever the body is in motion, u use F = uN.....but w.r.t the question u need to understand which friction is in play - static or kinetic....as ujjwal pointed out.

OpenStudy (lgbasallote):

wait..\[F_N = F_g??\]

OpenStudy (anonymous):

nooo...Fn = normal force

OpenStudy (ujjwal):

In this case \(F_N=mg\)

OpenStudy (vincent-lyon.fr):

@ujjwal quote "Once the body is in motion, F=μsN" This is not correct. Here, the tire is not slipping, so F<µsN but, as we want the most effective braking, we will brake so that F is maximum, so F will be nearly equal to µsN. This is why the equations will be solved using µsN = µs mg for the retarding force.

OpenStudy (ujjwal):

@Vincent-Lyon.Fr Isn't rolling friction equal to static friction?

OpenStudy (vincent-lyon.fr):

Main friction, when rolling is static. "Rolling friction" is something ELSE that is always neglected in such problems.

OpenStudy (vincent-lyon.fr):

I mean your remark is ok, but you should just avoid using the phrase "rolling friction" which is incorrect.

OpenStudy (lgbasallote):

okay...i get the acceleration...explanation for the second equation?

OpenStudy (ujjwal):

@Vincent-Lyon.Fr When the body is rolling, the friction which comes into play is static friction but then you say F<\(\mu\)s N.. Shouldn't it be F=\(\mu\)s N

OpenStudy (anonymous):

@Igbasallote then u just have to substitute value of a in v^2 = u^2 + 2as.....like how ujjwal showed above.

OpenStudy (lgbasallote):

lol people need to stop tagging igbasallote..it's LGBASALLOTE

OpenStudy (anonymous):

ohhh.. now wonder ur name was not showing up!

OpenStudy (vincent-lyon.fr):

@ujjwal it is F≤μs N in the general case, and F=μs N in this example, because we are looking for the shortest braking distance, which implies the maximum braking force.

OpenStudy (ujjwal):

I guess i need to revise my concepts on friction.. Thanks @Vincent-Lyon.Fr ..

OpenStudy (lgbasallote):

that second equation is similar to \[V^2 = V_o^2 + 2a(x_2 - x_1)\] right?

OpenStudy (ujjwal):

yes!

OpenStudy (lgbasallote):

why 15^2 shouldnt it be 50^2?

OpenStudy (anonymous):

right...that should be 50^2

OpenStudy (ujjwal):

yeah it should be 50^2.. sorry, my mistake.

OpenStudy (lgbasallote):

okay i tink i got the process in this one..onto more practices

OpenStudy (ujjwal):

best of luck, buddy!

OpenStudy (anonymous):

gr8....work hard!